# Cell divisions imprint long lasting elastic strain fields in epithelial tissues

### v1

@@ Abstract
A hallmark of biological tissues, viewed as complex cellular materials, is the active generation of mechanical stresses by cellular processes, such as cell divisions. Each cellular event generates a force dipole that deforms the surrounding tissue. Therefore, a quantitative description of these force dipoles, and their consequences on tissue mechanics, is one of the central problems in understanding the overall tissue mechanics. In this work we analyze previously published experimental data on fruit fly D. melanogaster wing epithelia to quantitatively describe the deformation fields induced by a cell-scale force dipole. We find that the measured deformation field can be explained by a simple model of fly epithelium as a linearly elastic sheet. This fact allows us to use measurements of the strain field around cellular events, such as cell divisions, to infer the magnitude and dynamics of the mechanical forces they generate. In particular, we find that cell divisions exert a transient isotropic force dipole field, corresponding to the temporary localisation of the cell nucleus to the tissue surface during the division, and traceless-symmetric force dipole field that remains detectable from the tissue strain field for up to about $3.5$ hours after the division. This is the timescale on which elastic strains are erased by other mechanical processes and therefore it corresponds to the tissue fluidization timescale. In summary, we have developed a method to infer force dipoles induced by cell divisions, by observing the strain field in the surrounding tissues. Using this method we quantitatively characterize mechanical forces generated during a cell division, and their effects on the tissue mechanics.
@@ I Introduction
How mechanical forces influence biological tissues is one of the central problems in animal development and regeneration Heisenberg and Bellaïche (2013); Brugués et al. (2014); LeGoff and Lecuit (2016); Jülicher and Eaton (2017); Tetley et al. (2019). Biological tissues are often described as soft viscoelastic materials Aigouy et al. (2010) that are elastic on sufficiently short timescales. Tissue fluidization on longer timescales arises from cell divisions, extrusions and intercalations, that restructure the tissue over time Ranft et al. (2010); Etournay et al. (2015); Curran et al. (2017); Kim et al. (2021).
@@
Cell divisions are an intrinsic source of mechanical force generation within tissues. When a cell divides, particularly in a closely packed epithelium, it induces deformation in the surrounding tissue Gibson et al. (2011); Mao et al. (2013), potentially leading to cellular rearrangements Firmino et al. (2016) and contributing to the overall tissue flow Ranft et al. (2010); Matoz-Fernandez et al. (2017); Bocanegra-Moreno et al. (2023). Furthermore, cell divisions have been indicated to control the glassy cell dynamics observed in cultured epithelial tissues Angelini et al. (2011); Park et al. (2015) and explanted embryonic tissues Schotz et al. (2013). However, despite the fact that cell divisions are a hallmark of biological matter and seem to be major factor in controlling material properties of biological tissues, quantitative understanding of the mechanical forces generated by cell divisions and their immediate consequences on the surrounding tissue is still lacking.
@@
In order to study the mechanics of cell divisions, we need to probe mechanical stresses in the surrounding tissue. One of the most common methods of probing mechanical stresses is the laser ablation experiment, in which one or multiple cells are destroyed by a laser. The resulting displacement of the ablated tissue boundary reflects the pre-stresses existing in the tissue before the ablation Farhadifar et al. (2007); Mayer et al. (2010); Bonnet et al. (2012); Kasza et al. (2014); Etournay et al. (2015); Piscitello-Gómez et al. (2023). Information about the pre-stresses are extracted either from the initial dynamics of tissue boundary displacement Mayer et al. (2010); Kasza et al. (2014) or from the final relaxed shape of the tissue Farhadifar et al. (2007); Bonnet et al. (2012); Dye et al. (2021); Piscitello-Gómez et al. (2023). Previous analyses focused on changes in position or shape of specific points or outlines in the vicinity of the ablation. Therefore, they could not provide an insight into the strain field induced in the surrounding tissue, and they had to rely on assumptions about tissue mechanical properties. Furthermore, to extract information about all components of the pre-stress in the tissue, previous methods require a precise, circular laser ablation Bonnet et al. (2012); Dye et al. (2021). If we could instead infer force dipoles induced by cellular processes such as cell divisions, we would be able to extract information about tissue mechanics in a completely non-invasive way. Such a method was recently developed to infer the properties of the core of plastic particle rearrangements in computer glasses Moriel et al. (2020, 2024). Therefore, the starting point of our work, on the way to study cell division mechanics, is to develop a method that allows us infer the force dipole in an epithelial tissue from the tissue strain field.
@@
In this work, we first provide a detailed quantitative analysis of the tissue displacement field following a linear laser ablation performed in the pupal wing epithelium of the fruit fly D. melanogaster. We show that the response of the fly wing epithelium to a laser ablation is consistent with that of a two-dimensional linear elastic sheet. This allows us to infer the force dipole induced by the ablation, normalized by the elastic constant, from the observed strain field. Furthermore, by analyzing the dynamics of the strain field, we show that the dominant mode of dissipation in the fly wing epithelium on short timescales of $1s$ to $50s$ is the viscous dissipation, as opposed to frictional dissipation with a substrate.
@@
Using our method to infer force dipole from its strain field, we determine the force dipole generated by cell divisions in the D. melanogaster wing disc. We find that the isotropic component of the cell division force dipole tensor is only transient and vanishes almost immediately after the division, indicating that no net cell growth is accumulated during the cell division. Furthermore, we find that the traceless-symmetric component of the force dipole, imprinted by the cell division, remains visible in the surrounding tissue for up to about $3.5$ hours after the division. This provides a measurement of the fluidization timescale in the tissue, and demonstrates importance of elastic interactions in tissue-scale morphogenetic processes that take place on a comparable timescale, such as wing disc eversion Fuhrmann et al. (2023).
@@ II.1 Linear laser ablation experiments
Linear laser ablation experiments, during which a small linear segment of tissue is destroyed by a laser, have been performed in the D. melanogaster wing epithelium during pupal morphogenesis Piscitello-Gómez et al. (2023). To obtain information about the pre-stress in the tissue, the dynamics of the closest cell-cell junction was measured as it retracted away from the ablation. Here, we use movies of laser ablation experiments published in Ref. Piscitello-Gómez et al. (2023) to study the deformation in the tissue surrounding the laser ablation. An example of such a laser ablation experiment is shown in Fig. 1 (A.i), where the red line indicates the line ablated by the laser. Following the ablation, the tissue relaxes until reaching a mechanical equilibrium after at most $T_{e}=50s$ in Fig. 1 (A.ii), see Methods.
@@
To measure the tissue strain field, we segmented the apical cell surfaces over time and quantified the area $a$ and the elongation tensor $Q_{ij}$ of each cell, which quantifies its shape Etournay et al. (2015); Merkel et al. (2017). Since no cell divisions or rearrangements appear in the region of interest following the laser ablation, we quantify the tissue strain field ${U}_{ij}(\bm{r},t)$ from the change in cell shape and size relative to their values before the ablation, see Methods. The strain field tensor consists of the isotropic strain component $U_{kk}$ and traceless-symmetric component $\tilde{U}_{ij}$ which we denote as pure shear strain component. In Fig. 1 (A.iv) we show the pure shear strain field component $\tilde{U}_{xy}$ at $100s$ after the laser ablation, which exhibits a clear spatial pattern, see SI Fig. 2 (A) for the other traceless-symmetric component and isotropic component of experiment 1 and SI Fig. 3 (A) for all components of the experiment 2.
@@
Our goal is to find the relationships between the induced strain field and the forces induced by the ablation. Since the process of ablation does not introduce any net force, the dominant component of the induced force field will be dipolar. Therefore, we aim to relate components of the strain field at distance $\vec{r}$ from the ablation to the components of the force dipole tensor, as illustrated in Fig. 1 (B.i,ii).
@@
We first consider tissue strain after it has reached the mechanical equilibrium after the time $T_{e}$ and calculate the discrete angular Fourier spectrum $\tilde{U}^{ss}_{ij}(r,m)$, where and $r=|\bm{r}|$ is distance from the center of ablation and $m$ is the angular mode index, see Methods. We find the strongest signal in the second and the fourth mode of the pure shear strain components, see Fig. 1 (C.i) and SI Fig. 2 (B.ii), which correspond to $\cos(2\varphi)$ and $\cos(4\varphi)$ angular Fourier modes. In the isotropic strain component we find the strongest signal in zeroth and second mode, see SI Fig. 2 (B.i), which correspond to the constant and $\cos(2\varphi)$ angular Fourier modes. We then quantified the radial dependence of the second angular mode and we find that its magnitude decays as $r^{-2}$, as shown in Fig. 1 (C.ii). These results reminded us of the far field of Eshelby propagators in two-dimensional linear elastic sheets Eshelby (1957); Nicolas et al. (2018); Moriel et al. (2020) and we next tested whether linear elastic theory can indeed quantitatively account for our measurements.
@@
We briefly present linear elastic theory of a sheet in which a point force dipole has been inserted. A force dipole is represented by a tensor $D_{ij}$ that can be decomposed into isotropic and traceless-symmetric contributions $D_{ij}=D_{0}\delta_{ij}+\tilde{D}_{ij}$  [^1] We consider only force dipoles that do not exert a net torque.. Inserting a force dipole in an elastic sheet induces a strain field that is determined by the force balance equation
@@
where $G_{ijkl}$ is the elastic dipole propagator and $\bm{r}$ is the distance from the force dipole, see SI. We graphically illustrate the components of Eq. 2 in Fig. 1 (B.iii), using the Voigt tensor representation. In this representation, the force dipole propagator elements are organised in a matrix, which we plot in Fig. 1 (b). From this point, we use the normlised force dipole defined as $d_{ij}=D_{ij}/(2\pi K(1+\mu))$. The propagator elements depend only on $\cos(2\varphi)$, $\sin(2\varphi)$, $\cos(4\varphi)$ and $\sin(4\varphi)$, see Fig. 1 (B.iii) and SI, consistent with the observed second and fourth angular Fourier modes in the data. Furthermore, all components of $G_{ijkl}$ decay in space as $r^{-2}$. This is consistent with the the radial profile observed in the experimental data for $m=2$ modes, see Fig. 1 (C.ii) and SI Fig. 4. Radial profile of $m=4$ modes decay faster and appear to be much more susceptible to the noise in the data. Therefore, the strain field measured in the laser ablation experiments is consistent with a linear elastic response of the tissue to the force dipoles generated by the ablation. However, to confirm this hypothesis, we need to show that linear elastic theory can quantitatively account for the observed strain fields.
@@
We fit the steady state tissue strain field $U^{ss}_{ij}$ with a strain field predicted by a linear elastic theory of a two-dimensional elastic sheet in Eq. 2. The fitting parameters are the three independent components of the force dipole tensor and the ratio $\mu$ of the two-dimensional shear and bulk elastic moduli. By analysing two laser ablation experiments, we find that the data are well described by the fit, as shown in Fig. 1 (D) and SI Fig. 5. The fitting parameters are consistent between the two experiments, see Table 1, except $\tilde{d}_{xx}$, which varies by a factor of two. This shows that there is a difference in the force dipole magnitude generated by the ablation, which could arise either from a difference of tissue pre-stresses or a difference in the effective ablation size in the two experiments. Furthermore, the ratio $\mu\approx 2$, corresponding to the Poisson ratio of $\nu\approx-0.33$, is comparable with the value $\mu\approx 3$, corresponding to the Poisson ration of $\nu\approx-0.5$, recently reported in the same tissue using a laser ablation method based on circular ablation shape Piscitello-Gómez et al. (2023).
@@
To further test our analysis of the laser ablation experiments, we simulate a linear ablation in the vertex model of epithelial tissues Farhadifar et al. (2007). We remove two cells from a pre-stressed cellular network, see SI, and then we quantify the resulting strain. We find that the simulation shows the same pattern as in experiments, consistent with the linear elastic theory of a 2-d sheet Fig. 1 (E, F) and SI Fig. 6. We fit the linear elastic theory to the observed strain field pattern and infer the force density generated by the ablation from tissue deformation. Finally, we find the Poisson’s ratio of the vertex model cellular network to be $\nu=-0.061\pm 0.005$. This finding is consistent with the value predicted value $\nu\approx-0.060$ in Ref. Staddon et al. (2023) for the vertex model parameters we use, see SI.
@@
Our results show that response of the D. melanogaster wing epithelium to a small linear laser ablation is consistent with that of a linearly elastic sheet with a negative two-dimensional Poisson ratio. This allows us to infer the force dipole tensor induced by the ablation from the tissue strain field, which we confirmed using the vertex model simulations.
@@
Figure 1: (A.i) Fluorescently labeled cell outlines show the cellular packing before the laser ablation. The red bar indicates the location that will be ablated. (A.ii) Following the laser ablation tissue has deformed away from the ablated region. (A.iii) The cells we have tracked and segmented define the region of interest colored in white. (A.iv) In the region of interest, we show the accumulated cell strain component $U_{xy}$. The scale bar is $10\mu m$. (B.i) A force dipole in a two dimensional linear elastic system induces a strain in distance $\vec{r}$. (B.ii) Two dimensional strain field and force dipole tensors are decomposed into isotropic and traceless-symmetric components. (B.iii) A Voigt representation of the relation between strain at position $\vec{r}$ induced by a force dipole at the origin. (C.i)Angular Fourier modes of $\tilde{U}_{xy}$ at distance $r=16~{}\rm{\mu m}$, which corresponds to the yellow circle A.iv, shows a clear signal in modes $m=2$ and $m=4$ modes. (C.ii) The radial profile of the $m=2$ angular mode shows a radial profile consistent with $r^{-2}$ (dashed line). The dotted line indicates the distance $r=16~{}\rm{\mu m}$. (D.i) Fit of the linear elastic theory to the experimental data. The solid blue line is the fitted strain field with the corresponding fit parameters reported in Table I, exp. 1. (D.ii) Visualisation of the fitted strain field. (E.i) The cellular network of tissue in the vertex model simulation. The red bar indicates two cells removed by an ablation. (E.ii) The cell strain field accumulated in the tissue after the ablation. (F.i) Angular Fourier modes of the strain field pattern in E.ii showing peaks in the second and fourth modes. (F.ii) Fit of the linear elastic theory to cell strain field in the vertex model. The solid blue is the fitted strain field with the fit parameters $d_{0}=(10.92\pm 0.01)\times 10^{-2}a_{0}$, $\tilde{d}_{xx}=(8.29\pm 0.01)\times 10^{-2}a_{0}$, $\tilde{d}_{xy}=(2.93\pm 0.01)\times 10^{-2}a_{0}$ and $\mu=1.13\pm 0.01$.

### v2

@@ Abstract
A hallmark of biological tissues, viewed as complex cellular materials, is the active generation of mechanical stresses by cellular processes, such as cell divisions. Each cellular event generates a force dipole that deforms the surrounding tissue. Therefore, a quantitative description of these force dipoles, and their consequences on tissue mechanics, is one of the central problems in understanding the overall tissue mechanics. In this work, we analyze previously published experimental data on fruit fly D. melanogaster wing epithelia to quantitatively describe the deformation fields induced by a cell-scale force dipole. We find that the measured deformation field can be explained by a simple model of wing epithelium as a linearly elastic sheet. This allows us to infer the magnitude and dynamics of the mechanical forces generated by the cell divisions. In particular, we find that cell divisions exert a transient isotropic force dipole field, corresponding to the temporary localization of the cell nucleus to the tissue surface during the division, and traceless-symmetric force dipole field that remains detectable from the tissue strain field for up to about $3.5$ hours after the division. This is the timescale on which elastic strains are erased by other mechanical processes and therefore it corresponds to the tissue fluidization timescale. In summary, we have developed a method to infer force dipoles induced by cell divisions, by observing the strain field in the surrounding tissues. Using this method we quantitatively characterize mechanical forces generated during a cell division and their effects on the tissue mechanics.
@@ I Introduction
How mechanical forces influence biological tissues is one of the central problems in animal development and regeneration Heisenberg and Bellaïche (2013); Brugués et al. (2014); LeGoff and Lecuit (2016); Jülicher and Eaton (2017); Tetley et al. (2019). Biological tissues are often described as soft viscoelastic materials Aigouy et al. (2010) that are elastic on sufficiently short timescales. Tissue fluidization on longer timescales arises from cell divisions, extrusions, and intercalations, which restructure the tissue over time Ranft et al. (2010); Etournay et al. (2015); Curran et al. (2017); Kim et al. (2021).
@@
Cellular processes, such as cell divisions and cell extrusions, are intrinsic sources of mechanical force generation within tissues, inducing deformation in the surrounding tissue Gibson et al. (2011); Mao et al. (2013); Roellig et al. (2022), which can in turn lead to cellular rearrangements Firmino et al. (2016) and contribute to the overall tissue flow Ranft et al. (2010); Matoz-Fernandez et al. (2017); Bocanegra-Moreno et al. (2023). Furthermore, cell divisions have been indicated to control the glassy cell dynamics observed in cultured epithelial tissues Angelini et al. (2011); Park et al. (2015) and explanted embryonic tissues Schotz et al. (2013). However, despite the fact that cell divisions are a hallmark of biological matter and seem to be a major factor in controlling material properties of biological tissues, a quantitative understanding of the mechanical forces generated by cell divisions and their immediate consequences on the surrounding tissue is still lacking.
@@
To study the mechanics of cell divisions, we need to probe mechanical stresses in the surrounding tissue. One of the most common methods of probing mechanical stresses is the laser ablation experiment, in which one or multiple cells are destroyed by a laser. The resulting displacement of the ablated tissue boundary reflects the pre-stresses existing in the tissue before the ablation Farhadifar et al. (2007); Mayer et al. (2010); Bonnet et al. (2012); Kasza et al. (2014); Etournay et al. (2015); Piscitello-Gómez et al. (2023). Pre-stresses are then characterised either by the initial dynamics of tissue boundary displacement Mayer et al. (2010); Kasza et al. (2014) or from the final relaxed shape of the tissue Farhadifar et al. (2007); Bonnet et al. (2012); Dye et al. (2021); Piscitello-Gómez et al. (2023). Previous analyses focused on changes in the position or shape of specific points or outlines in the vicinity of the ablation. Therefore, they could not provide an insight into the strain field induced in the surrounding tissue, and they had to rely on assumptions about tissue mechanical properties. Furthermore, to extract information about all components of the pre-stress in the tissue, previous methods require a precise, circular laser ablation Bonnet et al. (2012); Dye et al. (2021). If we could instead infer force dipoles induced by cellular processes such as cell divisions, we would be able to extract information about tissue mechanics in a completely non-invasive way. Therefore, the starting point of our work, on the way to study cell division mechanics, is to develop a method that allows us to infer the force dipole in an epithelial tissue from the tissue strain field. The idea to infer properties of a localized force field from the strain field it induces was recently employed to characterize the properties of the core of plastic events in computer glasses Moriel et al. (2020); Chacko et al. (2021); Moriel et al. (2024). Here, we proceed independently of these results using an approach better suited for epithelial tissue data.
@@
In this work, we first provide a detailed quantitative analysis of the tissue displacement field following a linear laser ablation performed in the pupal wing epithelium of the fruit fly D. melanogaster. We show that the response of the fly wing epithelium to a laser ablation is consistent with that of a two-dimensional linear elastic sheet. This allows us to infer the force dipole induced by the ablation, normalized by the elastic constant, from the observed strain field. Furthermore, by analyzing the dynamics of the strain field, we show that the dominant mode of dissipation in the fly wing epithelium on short timescales less than a minute is viscous dissipation, as opposed to frictional dissipation with a substrate.
@@
Using our method to infer force dipole from its strain field, we determine the force dipole generated by cell divisions in the D. melanogaster wing disc. We find that the isotropic component of the cell division force dipole tensor is only transient and vanishes almost immediately after the division, indicating that no net cell growth is accumulated during the cell division. Furthermore, we find that the traceless-symmetric component of the force dipole, imprinted by the cell division, remains visible in the surrounding tissue for up to about $3.5$ hours after the division. This provides a measurement of the fluidization timescale in the tissue and demonstrates the importance of elastic interactions in tissue-scale morphogenetic processes that take place on a comparable timescale, such as wing disc eversion Fuhrmann et al. (2023).
@@ II.1 Linear laser ablation experiments
In linear laser ablation experiments, a small linear segment of tissue is destroyed by a laser. In Ref. Piscitello-Gómez et al. (2023), laser ablation experiments have been performed in the D. melanogaster wing epithelium during pupal morphogenesis. Information about the pre-stresses in the tissue was inferred from the dynamics of the nearest cell-cell junction as it retracted away from the ablation. Here, we use laser ablation experiment data published in Ref. Piscitello-Gómez et al. (2023) to study the deformation field in the tissue surrounding the laser ablation. An image of a wing tissue before and after such a laser ablation experiment is shown in Fig. 1 (A.i), where the red line indicates the line ablated by the laser. Ablations were performed in a plane thinner than $1~{}\mu m$ on the apical surface of the tissue Piscitello-Gómez et al. (2023). Following the ablation, the tissue relaxes until reaching a mechanical equilibrium after about $T_{e}\approx 30s$ after ablation. We show an example of the relaxed state in Fig. 1 (A.ii). The tissue remains in this state until after around $1\rm{min}$. After this time, the tissue starts to exhibit additional flows, directed towards the line of ablation, which likely correspond to the wound-healing response of the tissue, see Methods and SI. Note that in most experiments from Ref. Piscitello-Gómez et al. (2023) we could not segment and analyze a reasonably large circle of cells around the ablation (SI Fig. 1), we selected four experiments that we could best analyze. In the main text we present analysis of the clearest one, denoted Experiment 1. In Experiment 2, such flows appear earlier so that no intermediate relaxed state can be identified. In Experiments 3 and 4 the strain field generated by the ablation is also weaker than in Experiment 1, so that the noise overcomes the signal already at small distances from the ablation. Analyses of these three experiments are presented in the SI.
@@
To measure the tissue strain field, we segment the apical cell surfaces over time and quantify area $a$ and the elongation tensor $Q_{ij}$ of each cell. The cell elongation tensor is defined such that the co-rotational changes in the elongation tensor $Q_{ij}$ correspond to the accumulated pure shear strain Etournay et al. (2015); Merkel et al. (2017). Since no cell divisions or rearrangements appear in the region of interest following the laser ablation, we quantify the tissue strain field ${U}_{ij}(\bm{r},t)$ from the change in cell shape and size relative to their values before the ablation, see Methods. The strain field tensor consists of the isotropic strain component $U_{kk}$ and traceless-symmetric component $\tilde{U}_{ij}$ which we denote as pure shear strain component. In Fig. 1 (A.iv) we show the pure shear strain field component $\tilde{U}_{xy}$ at $50s$ after the laser ablation, which exhibits a clear spatial pattern. See SI Fig. 3 (A) for the other traceless-symmetric component and the isotropic component.
@@
Our goal is to find the relationships between the induced strain field and the stresses induced by the ablation. Since the process of ablation does not introduce any net force, the dominant component of the induced force field will be dipolar. Therefore, we aim to relate components of the strain field at distance $\vec{r}$ from the ablation to the components of the force dipole tensor, as illustrated in Fig. 1 (B.i,ii).
@@
We first consider tissue strain after it has reached the mechanical equilibrium after the time $T_{e}$ and calculate the discrete angular Fourier spectrum $\tilde{U}^{ss}_{ij}(r,m)$, where $r=|\bm{r}|$ is the distance from the center of ablation and $m$ is the angular mode index, see Methods. We find the strongest signal in the second and the fourth mode of the pure shear strain components, see Fig. 1 (C.i) and SI Fig. 3 (B.ii), which correspond to $\cos(2\varphi)$ and $\cos(4\varphi)$ angular Fourier modes. In the isotropic strain component, we find the strongest signal in the zeroth and second mode, see SI Fig. 3 (B.i), which correspond to the constant and $\cos(2\varphi)$ angular Fourier modes. We then quantify the radial dependence of the second angular mode and we find that its radial profile is consistent with $r^{-2}$, as shown in Fig. 1 (C.ii). These results reminded us of the far field of Eshelby propagators in two-dimensional linear elastic sheets Eshelby (1957); Nicolas et al. (2018); Moriel et al. (2020), and we next tested whether the linear elastic theory can indeed quantitatively account for our measurements.
@@
We briefly present the linear elastic theory of a sheet in which a point force dipole has been inserted. A force dipole is represented by a tensor $D_{ij}$ that can be decomposed into isotropic and traceless-symmetric contributions $D_{ij}=D_{0}\delta_{ij}+\tilde{D}_{ij}$  [^1] We consider only force dipoles that do not exert a net torque.. Inserting a force dipole in an elastic sheet induces a strain field that is determined by the force balance equation
@@
where $G_{ijkl}$ is the elastic dipole propagator and $\bm{r}$ is the distance from the force dipole, see SI. We graphically illustrate the components of Eq. 2 in Fig. 1 (B.iii), using the Voigt tensor representation. In this representation, the force dipole propagator elements are organized in a matrix, which we plot in Fig. 1 (B). From this point, we use the normalized force dipole defined as $d_{ij}=D_{ij}/2K$. The propagator elements depend only on $\cos(2\varphi)$, $\sin(2\varphi)$, $\cos(4\varphi)$ and $\sin(4\varphi)$, see Fig. 1 (B.iii) and SI, consistent with the observed second and fourth angular Fourier modes in the data. Furthermore, all components of $G_{ijkl}$ decay in space as $r^{-2}$. This is consistent with the the radial profile observed in the experimental data for $m=2$ modes, see Fig. 1 (C.ii) and SI Fig. 5. The radial profiles of $m=4$ modes are more noisy so that it is difficult to characterise the radial decay.
Therefore, the strain field measured in the laser ablation experiments is consistent with a linear elastic response of the tissue to the force dipoles generated by the ablation. However, to confirm this hypothesis, we need to show that linear elastic theory can quantitatively account for the observed strain fields.
@@
We fit the steady state tissue strain field $U^{ss}_{ij}$ with a strain field predicted by a linear elastic theory of a two-dimensional elastic sheet in Eq. 2. The fitting parameters are the three independent components of the force dipole tensor and the ratio $\mu$ of the two-dimensional shear and bulk elastic moduli. By analyzing the ablation experiment, we find that the data are well described by the fit, as shown in Fig. 1 (D) and SI Fig. 6 (A). The fitted parameter values are: $d_{0}=160\pm 10~{}\mu m^{2}$, $\tilde{d}_{xx}=150\pm 10~{}\mu m^{2}$, $\tilde{d}_{xy}=20\pm 10~{}\mu m^{2}$, $\mu=1.9\pm 0.1$, see Methods. Furthermore, the ratio $\mu\approx 2$, corresponding to the Poisson ratio of $\nu\approx-0.33$, is comparable with the value $\mu\approx 3$, corresponding to the Poisson ratio of $\nu\approx-0.5$, recently reported in the same tissue using a laser ablation method based on circular ablation shape Piscitello-Gómez et al. (2023).
@@
To further test our analysis of the laser ablation experiments, we simulate a linear ablation in the vertex model of epithelial tissues Farhadifar et al. (2007). We remove two cells from a pre-stressed cellular network, where the pre-stress represents the proximal-distal stress in the wing epithelium Etournay et al. (2015), and the ablation axis is perpendicular to the pre-stress orientation, recapitulating the orientation of ablation in experiments, see SI for details. We then quantify the resulting strain and find that the simulation shows the same pattern as in experiments, consistent with the linear elastic theory of a 2D sheet Fig. 1 (E, F) and SI Fig. 7. We fit the linear elastic theory to the observed strain field pattern and infer the force density generated by the ablation from tissue deformation. Finally, we find the Poisson’s ratio of the vertex model cellular network to be $\nu=-0.13\pm 0.04$. This finding is consistent with the value $\nu=-0.095\pm 0.009$ found by a direct numerical measurement of the shear and bulk modulus of the same network, see SI.
@@
Our results show that the response of the D. melanogaster wing epithelium to a small linear laser ablation is consistent with that of a linearly elastic sheet with a negative two-dimensional Poisson ratio. This allows us to infer the force dipole tensor induced by the ablation from the tissue strain field, which we confirmed using the vertex model simulations.
@@
Figure 1: (A.i) Fluorescently labeled cell outlines show the cellular packing before the laser ablation. The red bar indicates the location that will be ablated. (A.ii) Cell outlines $50s$ after the laser ablation. (A.iii) Tracked and segmented cells in the region of interest are colored in white. (A.iv) The accumulated cell strain component $U_{xy}$, $50s$ after the laser ablation. The scale bar is $10\mu m$. (B.i) A force dipole in a two-dimensional linear elastic system induces a strain in distance $\vec{r}$. (B.ii) Decomposition of two-dimensional strain field and force dipole tensors into isotropic and traceless-symmetric components. (B.iii) A Voigt representation of the relation between strain at position $\vec{r}$ induced by a force dipole at the origin. (C.i) Angular Fourier modes of $\tilde{U}_{xy}$ at distance $r=16~{}\rm{\mu m}$, which corresponds to the yellow circle A.iv, show a clear signal in modes $m=2$ and $m=4$. (C.ii) The radial profile of the $m=2$ angular mode shows a radial profile consistent with $r^{-2}$ (dashed line). The dotted line indicates the distance $r=16~{}\rm{\mu m}$. Analysis in both C.i and C.ii is performed on the strain field averaged over the time interval from $30s$ to $70s$ after laser ablation and the error bars show the standard deviation over this interval. (D.i) The fit of the linear elastic theory to the experimental data, shown in C. The solid blue line is the fitted strain field with the corresponding fit parameters reported in the text. (D.ii) Visualization of the fitted strain field. (E.i) The cellular network in the vertex model simulation. The red bar indicates two cells removed by ablation. (E.ii) The cell strain field accumulated in the tissue after reaching the steady state following the simulated ablation. (F.i) Angular Fourier modes of the strain field pattern in E.ii showing peaks in the second and fourth modes. (F.ii) The fit of the linear elastic theory to the cell strain field in the vertex model. The solid blue is the fitted strain field with the fit parameters $d_{0}=0.72\pm 0.07a_{0}$, $\tilde{d}_{xx}=0.65\pm 0.07a_{0}$, $\tilde{d}_{xy}=0.00\pm 0.07a_{0}$ and $\mu=1.3\pm 0.1$. The uncertainty intervals for fit parameters are obtained by bootstrapping the data (see Methods).